OpenStudy (moonlight93):

a solution of 1.15g of solute dissolved in 25.0 mL of H2O at 25 degree C has a boiling point of 100.85 degree C. What is the molar mass of the solute if it is a nonvolatile nonelectrolyte and the solution behaves ideally ( s of H2O at 25 degree C = 0.997 g/mL) ? I need help with this problem! what formulas should I use, and how can I start with this? Thank you for any help.

OpenStudy (surry99):

Look up boiling point elevation

OpenStudy (surry99):

you will find a formula that will get you started.

OpenStudy (moonlight93):

can you check my work, and my final answer please. (100.85 - 100.00) °C / (0.512 °C/m) = 1.660 m (1.660 mol/kg) x (25.0 mL x (0.997 g/mL) / (1000 g/kg)) = 0.04138 mol (1.15 g) / (0.04138 mol) = 25.27 g/mol

OpenStudy (moonlight93):

@surry99

OpenStudy (surry99):

Method looks fine, math looks good as well.

OpenStudy (moonlight93):

thank you

OpenStudy (moonlight93):

the final answer should be 28